# Massive Particle Motion Around Horndeski Black Holes

### v1

@@ II Four-Dimensional Horndeski Black Hole
In this section, after reviewing the Horndeski theory, we will discuss a particular hairy black hole solution Rinaldi:2012vy of the Horndeski theory generated by a scalar field non-minimally coupled to the Einstein tensor. The action of the Horndeski theory Horndeski:1974wa is given by
@@ IV.1.2 Planetary orbits
where the index $P$ is for planetary orbits and Weierstrass-$\wp$ corresponds to the limit-fromWeierstrass-$\wp$-℘ -Weierstrass elliptic function with the Weierstrass invariants
@@
Perihelion precession. To compute the perihelion shift in a simple and transparent way, we first establish a connection with Kepler’s classical problem. Starting from Eq. (20), and performing the standard change of variables $u=1/r$, we obtain
@@
Planet Observed shift.    Seconds per century Latus rectum ($\ell$) (×$\times 10^{10}\,m$) Revolution frequency ($f_{rev}$). Orbits per annum $\qquad\gamma$ (×$\times 10^{6}\,m$)
@@ IV.1.3 Second kind trajectories
The particles follow second-kind trajectories when they originate at some point in spacetime and subsequently plunge into the horizon; these orbits appear when the energy lies at $10<E<$ 1. It is important to note that within this energy range and for different values of the angular momentum, multiple solutions may exist. These solutions can be expressed in terms of elementary functions or elliptic functions, depending on the specific configuration.
@@
[…]
@@ IV.2.1 Scattering trajectories
where the index $s$ is for the scattering orbits and Weierstrass-$\wp$ corresponds to the limit-fromWeierstrass-$\wp$-℘ -Weierstrass elliptic function with the Weierstrass invariants
@@ V Motion with $L=0$
$\tau(r)$ $=$ ±$\pm\int^{r}_{r_{0}}{r^{\prime}\,dr^{\prime}\over\sqrt{p_{2}(r^{\prime})}}\,,$ (123)
$t(r)$ $=$ ±$\pm\int^{r}_{r_{0}}{r^{\prime 3}\,dr^{\prime}\over(r^{\prime}-r_{+})(r^{\prime}-\rho_{2})\sqrt{p_{2}(r^{\prime})}}\,,$ (124)
@@ V.2 Unbounded trajectories: $E\geq 1$
$t\left(r\right)=\pm E\tau\left(r\right)\pm{E\over\sqrt{E^{2}-1}}\left(2MH(r)+\right.\,$ (145)
@@ A.1 Critical points
Hence, we can conclude that in the range $L\in[0,L_{C})$, there are no circular orbits, since we have two complex radial roots from the negativity of the discriminant (157) and one negative real radial root by the Descartes rule of signs on the polynomial equation (150), this last negative root remains in the other two next cases, so we will ignore it. When $L=L_{C}$, the discriminant is zero, so we obtain two equal positive real roots $r_{U}=r_{S}=r_{\text{ISCO}}$ that correspond to the innermost stable circular orbit. In the range $L\in(L_{C},\infty)$, the discriminant is positive, so there are two different positive real roots, then there are two circular orbits (unstable and stable). Furthermore, by parity symmetry on the polynomial (158), when $L$ is negative, the behavior is analogous by considering the negative real root $-L_{C}$.
@@ A.4 Analysis
Combining the three analyses above, we find that $L_{I}\leq L_{C}$, since the discriminants $\triangle_{c}(L)$ and $\triangle_{I}(L)$ imply that the polynomials in Eqs. (158) and (192) satisfy the inequality $P_{c}(L)\leq P_{I}(L)$. This holds because the difference $P_{c}(L)-P_{I}(L)$ is explicitly negative. Therefore, evaluating the polynomial inequality at $L=L_{C}$ yields the corresponding inequality for the angular momentum. Similarly, it follows that $L_{C}\leq L_{S}$. In general, for positive angular momentum, the inequality $L_{I}\leq L_{C}\leq L_{S}$ holds, while for negative angular momentum, the relation $-L_{S}\leq-L_{C}\leq-L_{I}$ is satisfied.

### v2

@@ II Four-Dimensional Horndeski Black Hole
In this section, after reviewing the Horndeski theory, we will discuss a particular hairy black hole solution Babichev:2017guv of the Horndeski theory generated by a scalar field non-minimally coupled to the Einstein tensor. The action of the Horndeski theory Horndeski:1974wa is given by
@@ IV.1.2 Planetary orbits
where the index $P$ is for planetary orbits and $\wp$ corresponds to the $\wp$-Weierstrass elliptic function with the Weierstrass invariants
@@
Perihelion precession. It should be noted that, unlike in General Relativity, a general Birkhoff’s theorem does not necessarily hold in Horndeski gravity. Therefore, in this work we explicitly assume that the vacuum solution employed provides an accurate description of the exterior gravitational field of a spherically symmetric source such as the Sun, in analogy with the Schwarzschild solution in GR Chandrasekhar:579245 . This assumption is expected to remain valid in the weak-field regime, where deviations from GR are small. So, to compute the perihelion shift in a simple and transparent way, we first establish a connection with Kepler’s classical problem. Starting from Eq. (20), and performing the standard change of variables $u=1/r$, we obtain
@@
Planet Observed shift.  Seconds per century Latus rectum ($\ell$) ($\times 10^{10}\,m$) Revolution frequency ($f_{rev}$). Orbits per annum $\qquad\gamma$ ($\times 10^{6}\,m$)
@@ IV.1.3 Second kind trajectories
The particles follow second-kind trajectories when they originate at some point in spacetime and subsequently plunge into the horizon; these orbits appear when the energy lies at $0<E<1$. It is important to note that within this energy range and for different values of the angular momentum, multiple solutions may exist. These solutions can be expressed in terms of elementary functions or elliptic functions, depending on the specific configuration.
@@
Second kind trajectories for $L>L_{C}$, and $E_{S}<E<E_{U}$. For these cases, we use the same solution from (62), but we replace the initial condition $r_{0}=r_{a}$ in the function $w(r)=1/(r_{2}-r)$ with $r_{2}=r_{f}$, and denote the new solution as $r_{\mathrm{SK}}(\phi)$, where (SK) refers to second-kind orbits, see Figure 7(b).
@@ IV.2.1 Scattering trajectories
where the index $s$ is for the scattering orbits and $\wp$ corresponds to the $\wp$-Weierstrass elliptic function with the Weierstrass invariants
@@ V Motion with $L=0$
$\tau(r)$ $=$ $\pm\int^{r}_{r_{0}}{r^{\prime}\,dr^{\prime}\over\sqrt{p_{2}(r^{\prime})}}\,$, (123)
$t(r)$ $=$ $\pm\int^{r}_{r_{0}}{r^{\prime 3}\,dr^{\prime}\over(r^{\prime}-r_{+})(r^{\prime}-\rho_{2})\sqrt{p_{2}(r^{\prime})}}\,$, (124)
@@ V.2 Unbounded trajectories: $E\geq 1$
$t\left(r\right)=\pm E\tau\left(r\right)\pm{E\over\sqrt{E^{2}-1}}\left(2MH(r)+\right$.\,
@@ VI Final remarks
Acknowlegements
We thank the anonymous referee for valuable comments and suggestions.
@@ A.1 Critical points
Hence, we can conclude that in the range $L\in[0,L_{C})$, there are no circular orbits, since we have two complex radial roots from the negativity of the discriminant (157) and one negative real radial root by the Descartes rule of signs on the polynomial equation (150), this last negative root remains in the other two next cases, so we will ignore it. When $L=L_{C}$, the discriminant is zero, so we obtain two equal positive real roots $r_{U}=r_{S}=r_{\text{ISCO}}$ that correspond to the innermost stable circular orbit. In the range $L\in(L_{C},\infty)$, the discriminant is positive, so there are two different positive real roots, then there are two circular orbits (unstable and stable). Furthermore, by parity symmetry on the polynomial (158), when $L$ is negative, the behavior is analogous by considering the negative real root $L_{C}$.
@@ A.4 Analysis
Combining the three analyses above, we find that $L_{I}\leq L_{C}$, since the discriminants $\triangle_{c}(L)$ and $\triangle_{I}(L)$ imply that the polynomials in Eqs. (158) and (192) satisfy the inequality $P_{c}(L)\leq P_{I}(L)$. This holds because the difference $P_{c}(L)-P_{I}(L)$ is explicitly negative. Therefore, evaluating the polynomial inequality at $L=L_{C}$ yields the corresponding inequality for the angular momentum. Similarly, it follows that $L_{C}\leq L_{S}$. In general, for positive angular momentum, the inequality $L_{I}\leq L_{C}\leq L_{S}$ holds, while for negative angular momentum, the relation $L_{S}\leq-L_{C}\leq-L_{I}$ is satisfied.